Operations Report – February 16th

Crew 205 Operations Report 16 – FEB – 2019

SOL: 6

Name of person filing report: Dave Masaitis

Non-nominal systems: None

Notes on non-nominal systems: N/A

Generator (hours run): Turned off at 0800 hrs (16th Feb),Turned on at 1530 hrs (16th Feb),

Solar— SOC % – Turned off Gen (16th Feb) 74%, Turned on Gen (16th Feb) 91%

Diesel Reading – 85%

Propane Reading – ~ 44%

Ethanol Free Gasoline – 10 gallons

Water (Auxillary tank) – Not in use

Water (Static tank) – around 60%, ~ 330 gallons

Auxillary to Static tank transfer – No

Gallons transferred: Not Applicable

Water in GreenHab – ~ 228.5 gallons

Water (loft) – 12 marked level

Static to Loft Pump used – yes

Water Meter: 01412700 units

Toilet tank emptied: No

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Yes

Hours: 50.4 hours

Beginning charge: 100%

Ending charge: 86%

Currently charging: Yes

Curiosity rover used: Not Used

Hours: 74.3 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Spirit rover used: Yes

Hours: 75.7

Beginning charge: 100%

Ending charge: 73%

Currently charging: Yes

Notes on rovers: N/A

ATV’s Used: N/A

Reason for use: N/A

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were Used today: 0

Notes on ATVs: None

HabCar used and why, where? N/A

Crew Car used and why, where? Not used

General notes and comments: No general notes.

Summary of internet: Nothing to report

Summary of suits and radios: Suits 1 and 6 had right-hand air tubes disconnect at the backpack during EVA. The helmet accompanying Suit 6 also had a broken zip-tie on the back panel of the helmet. Suits 3 and 4 are having intermittent fan problems despite having correct battery voltage. These will be checked and repaired this evening and tomorrow by D. Masaitis and V. Trivino.

One radio lost charge early on EVA, despite appearing to be at full charge prior to EVA. It was marked upon return, and will be thoroughly tested before being cleared for EVA use again.

Summary of Hab operations: The lightbulb in the rear (tunnel-side) airlock is out. Crew requires a replacement light bulb in order to repair. The handle hardware of the front airlock interior door was loose. Crew Engineer tightened 7/8 screws, but the top-left screw of the top attachment point would not thread at all. Replacement benchstock required to make a successful repair.

Summary of Green Hab operations: Nothing to report.

Summary of Science Dome operations: Hardware on the drawer set between the microscope table and laminar flow hood was excessively loose to where the drawer set was beinning to collapse. Repaired by tightening all of the hardware.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns and requests to Mission Support: Replacement lightbulb for rear airlock, replacement benchstock for front airlock door handle

Copyright © The Mars Society. All rights reserved. | Main Site